### v1

# Nickel and Diming Your GAN: A Dual-Method Approach to Enhancing GAN Efficiency via Knowledge Distillation

Abstract
In this paper, we address the challenge of compressing generative adversarial networks (GANs) for deployment in resource-constrained environments by proposing two novel methodologies: Distribution Matching for Efficient compression (DiME) and Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). DiME employs foundation models as embedding kernels for efficient distribution matching, leveraging maximum mean discrepancy to facilitate effective knowledge distillation. Simultaneously, NICKEL employs an interactive compression method that enhances the communication between the student generator and discriminator, achieving a balanced and stable compression process. Our comprehensive evaluation on the StyleGAN2 architecture with the FFHQ dataset shows the effectiveness of our approach, with NICKEL & DiME achieving FID scores of 10.45 and 15.93 at compression rates of 95.73% and 98.92%, respectively. Remarkably, our methods sustain generative quality even at an extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a large margin. These findings not only demonstrate our methodologies’ capacity to significantly lower GANs’ computational demands but also pave the way for deploying high-quality GAN models in settings with limited resources. Our code will be released soon.
1 Introduction
Generative Adversarial Networks (GANs) have attracted significant popularity as one of the most promising generative models, alongside the diffusion models [12, 53, 4, 47], in various computer vision tasks such as super-resolution [38, 28, 56], image editing [9, 51, 40], and image generation [20, 22, 18]. Particularly, thanks to their fast inference speed compared to diffusion models, GANs offer significant advantages for real-time applications[46, 15, 23]. However, despite their outstanding performance, the application of state-of-the-art GANs [5, 19, 20, 22, 49, 50, 18] on edge devices is constrained by their huge resource consumption.
Although compression methods have been extensively studied for classification tasks [27, 10, 8, 39, 54], their naïve application to generative models often leads to significant performance degradation [52, 55]. As shown in Fig. 1, to distill the rich knowledge from the teacher to the student, simple label matching is performed in classification models, whereas high-dimensional output matching is required in generative models. Moreover, in GANs, achieving optimal performance requires a delicate balance between the generator and discriminator during adversarial training, which becomes more difficult between the pruned generator and discriminator (see Fig. 3(a)).
(a) KD for classifier
(b) KD for conditional GAN
(c) KD for unconditional GAN
Figure 1: Comparison of knowledge distillation methods. (a) In classification tasks, the instance matching of output labels between the teacher and student is performed. Output labels are in low-dimensional space. Ideally, the outputs of the student and teacher are the same. (b) In conditional generative tasks, the instance matching of output images between the teacher and student is performed. Output images are in high dimensional space. The outputs of the student and teacher are similar (in terms of structure or background). (c) In unconditional generative tasks, the distribution matching of output images between the teacher and student is performed. There is no necessity for each input to have the same output.
Recently, several GAN compression methods [31, 35, 55, 16, 13, 14, 57, 60, 32, 58] have been proposed, but compressing unconditional GAN remains challenging. This is because conditional GANs require instance matching [57] as the teacher and student strive for similar outputs in a manner akin to classification tasks, whereas unconditional GAN compression demands distribution matching [16] (Fig. 1). There exist a few unconditional GAN compression studies [55, 35, 16, 57], but they either still suffer from significant performance degradation [55, 35, 16, 57] or require additional costs such as manual labeling [35] and MCMC sampling [16].
To address these problems, we first propose the Distribution Matching for Efficient compression (DiME). Most GAN compression methods utilized the embedding space (e.g., perceptual [35, 14], frequency [60]) because directly matching high-dimensional output images leads to significant performance degradation. Similarly, we leverage the foundation models (i.e., DINO [2, 37], CLIP [43]) as embedding kernels, which have shown successful applications with strong embedding power on various tasks [34, 61, 26]. Furthermore, Santos et al. and Yeo et al. [48, 58] have shown that neural networks can be considered as characteristic kernels to map into Reproducing Kernel Hilbert Space (RKHS), where matching the extracted features of two distributions is equivalent to matching the original distributions as the maximum mean discrepancy (MMD) critic [6, 7, 33, 30, 48]. Additionally, we propose to utilize the global features of the teacher generator to reduce the sampling error. While we ideally hope for the matching of population distributions between the teacher generator ($G^{T}$) and the student generator ($G^{S}$) through knowledge distillation, in reality, there is a sampling error due to the matching between sample distributions. Since the distribution of the $G^{T}$ is fixed, according to the law of large numbers, we can obtain nearly error-free statistics by precomputing a large number of samples from $G^{T}$. We provide detailed discussion in Sec. 4.4.
Figure 2: A schematic overview of our method. Our method consists of (a) Distribution Matching for Efficient compression (DiME), (b) Network Interactive Compression via Knowledge Exchange and Learning (NICKEL), and (c) adversarial loss. (a) matches the outputs between the teacher generator ($G^{T}$) and the student generator ($G^{S}$) via the foundation model $\phi$ in the embedding space. (b) matches the intermediate features between the teacher generator and the student discriminator ($D^{S}$). (c) represents the adversarial loss between the student generator and both the teacher discriminator ($D^{T}$) and the student discriminator.
In addition to DiME, to exploit the characteristic of GAN that consists of a generator and a discriminator, we propose Network Interactive Compression via Knowledge Exchange and Learning (NICKEL). In GAN training, Lee et al. [29] has shown that the discriminator can provide more meaningful signals as feedback by learning the semantic knowledge of the generator. Inspired by Lee et al. , we not only distill knowledge directly between the generators (i.e.., DiME), but also distill knowledge from the more informative $G^{T}$ to the student discriminator ($D^{S}$) by transmitting knowledge between generators via the discriminator indirectly. By utilizing $G^{T}$, we obtain two distinct advantages, Firstly, from the onset of training, $D^{S}$ learns the rich semantic knowledge embedded within the $G^{T}$. Secondly, the $G^{T}$ provides a wealth of knowledge surpassing that of $G^{S}$. Furthermore, we observe that NICKEL enhances the stability of GAN compression (see Fig. 3). To the best of our knowledge, NICKEL is the first method that distills the knowledge from $G^{T}$ to $G^{S}$ via the feedback of $D^{S}$ for model compression.
Our experimental results show that DiME outperforms existing state-of-the-art compression methods through knowledge distillation between $G^{T}$ and $G^{S}$. By applying DiME to StyleGAN2, which has a baseline FID of 4.02, resulted in FID scores of 11.25 and 18.32 at compression rates of 95.87% and 98.92%, respectively. This compares favorably to the state-of-the-art method [16], which achieves FID scores of 14.01 and 22.23 at the same compression rate. This demonstrates the power of using foundation models as embedding kernels for knowledge distillation. In addition, by using NICKEL with DiME, we further enhance the FID scores to 10.45 and 15.93 with improved stability, setting a new standard in GAN compression performance. It is worth to note that we achieve a reasonable performance with the FID of 29.38 at the extreme compression rate of 99.69%, surpassing the previous state-of-the-art performance by a significant margin.
Our contributions can be summarized as follows:
  • We propose DiME, an effective distillation method for GANs that ensures the matching output distributions between $G^{T}$ and $G^{S}$ via employing foundation models as kernels for MMD loss (Sec. 3.1). DiME outperforms existing GAN compression methods, achieving state-of-the-art performance in GAN compression at all compression rates.
  • We propose NICKEL that further enhances the distillation capability by providing more meaningful feedback from $D^{S}$. We observe that NICKEL leads to the improvement of stability (Sec. 3.2).
  • With NICKEL & DiME, our final model further raises the bar of the state-of-the-art. We find that our method shows a stable convergence with competitive generative performance, even at the extremely high compression rates of 99.69% (Sec. 4).
(a) Logits of D on ITGC[16]
(b) Logits of D on NICKEL & DiME
(c) Convergence of FID
[…]
2 Related work
2.1 GAN Compression
Recently, various GAN compression methods [31, 55, 13, 35, 32, 57, 16, 60, 58, 14] have been explored because directly applying compression methods, which were originally proposed for discriminative tasks, to GANs has led to significant performance degradation. However, most GAN compression methods have been explored in conditional GAN settings [31, 32, 60, 14], which are unsuitable for distribution matching in unconditional GAN compression (see Fig. 1). Occasionally, to address this problem, GAN compression methods have been explored [35, 16, 57], proposing better embedding spaces or distance metrics between $G^{T}$ and $G^{S}$. Wang et al. [55] proposed the GAN slimming (GS), a unified optimization framework, and emphasized that naïve application of compression methods leads to significant performance degradation due to the notorious instability of GANs. Liu et al. [35] proposed the content-aware GAN compression (CAGC) method, which focuses on only the contents of interest (e.g., object, face) to distill the knowledge, but this method requires additional costs due to manual labeling of contents. Li et al. [32] proposed the generator-discriminator cooperative compression (GCC) to maintain the nash-equilibrium between $G^{S}$ and $D^{S}$, but nash-equilibrium still cannot be maintained, in complex settings. Kang et al. [16] proposed the information-theoretic GAN compression (ITGC) by maximizing the mutual information between $G^{T}$ and $G^{S}$. ITGC requires a lot of computational costs due to the energy-based model and MCMC sampling. Xu et al. [57] proposed the StyleKD that focuses on the mapping network to achieve consistent outputs between $G^{T}$ and $G^{S}$. However, StyleKD can only be applied to networks based on StyleGAN. To address these issues, we propose Distribution Matching for Efficient compression (DiME), which matches the distribution between $G^{T}$ and $G^{S}$ via foundation kernels.
2.2 Foundation models
The foundation models [44, 45, 47, 41, 42, 1, 43, 2, 37] such as GPT [41, 42, 1], CLIP [43], and DINO [2, 37] are large-scale pre-trained models and are getting a lot of attention recently. Foundation models, which acquire general knowledge through large-scale training, are driving remarkable advancements on downstream tasks [61, 24, 34, 26]. For example, CLIP is a contrastive vision-language model trained so that positive pairs are clustered while negative pairs are scattered on a large dataset of 400 million image-text pairs. DINO is trained through self-supervised learning with self-distillation, passing all crops for the student and only the global view for the teacher to encourage local-to-global correspondences. In this paper, we utilize the CLIP and DINO, which have an image encoder, as kernels for better embedding spaces.
2.3 Discriminator regularization
Generally, GAN compression methods are focused on the $G^{S}$, thus it is applied in the form of generator regularization for knowledge distillation. On the other hand, GCC [32] emphasized the importance of considering not only the generator but also the discriminator to maintain the Nash equilibrium state between the compressed generator and discriminator. Similar phenomena were observed by several studies [3, 60, 14]. To address this issue, GCC used the selective activation discriminator, which partially activates the channels of the discriminator by utilizing the capacity constraint to maintain the Nash equilibrium state. However, GCC still shows significant performance degradation due to instability. In GAN training, Lee et al. [29] proposed generator-guided discriminator regularization (GGDR). GGDR showed that the discriminator can learn the semantic knowledge from the generator and lead to performance improvement of the generator by providing more powerful adversarial loss as feedback. However, GGDR cannot inject meaningful knowledge of the generator into the discriminator in the early stage because the initial generator is close to being a randomly initialized generator. Inspired by GGDR, we propose the Network Interactive Compression via Knowledge Exchange and Learning (NICKEL), which distills the knowledge from $G^{T}$ to $D^{S}$ and encourages powerful feedback from $D^{S}$ to $G^{S}$.
3 Method
This section introduces our two novel methodologies, DiME and NICKEL. Fig. 2 shows the overview schematic of our approach.
3.1 Knowledge Distillation with Foundation Kernels MMD
Generally, knowledge distillation (KD) minimizes the distance $d_{kd}$ (e.g., wavelet loss [60]) between the outputs of $G^{T}$ and $G^{S}$, encouraging $G^{S}$ to mimic $G^{T}$. We can achieve more effective knowledge distillation by designing a better distance metric. In this paper, we propose Distribution Matching for Efficient compression (DiME), which matches the distributions between $G^{T}$ and $G^{S}$ in the space embedded by foundation kernels $\phi$ as distance $d_{kd}$:
$\mathcal{L}_{KD}=d_{kd}(G^{T}(z),G^{S}(z))=\mathbb{E}[||\phi(G^{T}(z)),\phi(G^{S}(z))||_{1}]$ (1)
This is equivalent to using the MMD critic [6, 7, 33, 30, 48], a statistical method that matches two distributions in RKHS, assuming that the foundation kernels $\phi$ are characteristic kernels [48, 58].
While Eq. 1 generally shows good performance, we observe the tremendous performance degradation of all baselines (i.e., GS, CAGC, ITGC, Eq. 1) when $G^{S}$ has extremely few parameters (see Fig. 3(c)). As shown in Fig. 3(a), the Nash equilibrium breaks down when $G^{S}$ has fewer parameters, which consequently leads to the performance degradation of adversarial loss. To improve the stability of KD loss in the early stage, we utilize the global features of $G^{T}$. The global features are computed by inferring over a multitude of images rather than batch images, enabling the calculation of popular distribution statistics. Utilizing the global features mitigates the sampling error induced by the batch size in KD, with detailed discussion included in Sec. 4.4.
3.2 Network Interactive Compression via Knowledge Exchange and Learning
GAN utilizes a discriminator, which is a learnable network as the loss during the training of the generator. The performance of the generator is heavily influenced by the quality of feedback provided by the discriminator. GGDR [29] showed that during GAN training, the discriminator can learn semantic knowledge from the generator. Subsequently, the discriminator provides better feedback to the generator, thus improving the performance of the generator.
Inspired by GGDR, we propose NICKEL, which distills knowledge from $G^{T}$ into $D^{S}$ to provide more powerful feedback to $G^{S}$. NICKEL has advantages over simply applying GGDR to $G^{S}$ for two reasons. First, GGDR may struggle to provide meaningful information when $G^{S}$ resembles a random network during early training, whereas NICKEL can distill rich information from $G^{T}$ from the outset. Second, in GAN compression, due to the smaller network structure of $G^{S}$, GGDR cannot provide knowledge as rich as $G^{T}$. Therefore, we propose fine-tuning $D^{S}$ via NICKEL to learn information from $G^{T}$. However, fine-tuning $D^{S}$ using the NICKEL loss alone is insufficient to fully leverage the information from the pre-trained discriminator. Therefore, for adversarial learning, both $D^{T}$ and $D^{S}$ are employed.
The loss function of NICKEL can be formulated as follows:
$\mathcal{L}_{\texttt{NICKEL}}=\sum_{i=1}^{L}d_{\texttt{NICKEL}}(G_{i}^{T}(z),f_{i}(D_{i}^{S}(G^{T}(z)))),$ (2)
where $D_{i}^{S}(G^{T}(z))$ and $G_{i}^{T}(z)$ represent the feature maps of the $i$-th layer of $D^{S}$ and $G^{T}$, respectively. $f_{i}$ is a linear transform to match the shape of feature maps. As Lee et al. [29] mentioned, the knowledge of the generator contains a lot of semantic information. Therefore, we utilize the wavelet loss [60] for $d_{\texttt{NICKEL}}$, which is good for matching semantic information.
3.3 Training objective
In summary, our training loss for GAN compression is formulated as:
$\mathcal{L}=\mathcal{L}_{adv}+\lambda_{dino}\cdot\mathcal{L}_{dino}+\lambda_{clip}\cdot\mathcal{L}_{clip}+\lambda_{\texttt{NICKEL}}\cdot\mathcal{L}_{\texttt{NICKEL}},$ (3)
where $\lambda_{dino}$ and $\lambda_{clip}$ are the weights for the knowledge distillation, which utilizes the dino embedding and clip embedding, respectively. $\lambda_{\texttt{NICKEL}}$ is the weight for NICKEL loss $\mathcal{L}_{\texttt{NICKEL}}$ in Eq. 2. $\mathcal{L}_{adv}$ is the adversarial loss, which is the min-max objective function that includes both $D^{T}$ and $D^{S}$. $\mathcal{L}_{dino}$ and $\mathcal{L}_{clip}$ are the knowledge distillation losses of DiME in Eq. 1.
4 Experiments
4.1 Setups
4.1.1 Implementations.
We use the 20, 15, and 10 for $\lambda_{dino}$, $\lambda_{clip}$, and $\lambda_{\texttt{NICKEL}}$, respectively. We utilize the pretrained weights for CLIP [^1] https://openaipublic.azureedge.net/clip/models/afeb0e10f9e5a86da6080e35cf09123aca3b358a0c3e3b6c78a7b63bc04b6762/RN50.pt and DINO [^2] https://github.com/OpenGVLab/CaFo?tab=readme-ov-file as kernels. We use the same pruned generator as CAGC. We match 64 by 64 channels in StyleGAN2 and 8 by 8 channels in SNGAN for NICKEL on LH, HL, and HH components via Haar wavelet. To obtain the global features, we conduct 20,000 model inferences with the batchsize of 256.
Table 1: Comparison of FID scores of our methods and state-of-the-art compression methods on SNGAN for CIFAR-10. We compare the performance of various compression methods on SNGAN at compression rates of 74.88% and 90.85%. Distribution Matching for Efficient compression (DiME), in Eq. 1, shows superior performance compared to other state-of-the-art compression methods. Furthermore, NICKEL & DiME, which combines NICEKL (Eq. 2) and DiME, shows significant performance improvements. Particularly, at a compression rate of 74.88%, our method shows performance comparable to the full model.
[…]
SNGAN CIFAR-10 Full model [36] 3.36B - 17.71
CAGC [35] 0.85B 74.88% 40.45
ITGC [16] 43.66
DiME 31.98
NICKEL & DiME 23.11
CAGC [35] 0.31B 90.85% 51.93
ITGC [16] 59.99
DiME 36.89
NICKEL & DiME 28.27
Table 2: Comparison of FID scores of our methods and state-of-the-art compression methods on StyleGAN2 for FFHQ and LSUN-CAT. We compare the performance of various compression methods on StyleGAN2 for FFHQ and LSUN-CAT datasets at various compression rates. Distribution Matching for Efficient compression (DiME), in Eq. 1, outperforms state-of-the-art GAN compression methods and shows a significant performance gap compared to state-of-the-art GAN compression methods at high compression rates. Our final method, combining NICKEL (Eq. 2) with DiME, significantly enhances the performance. Even NICKEL & DiME, which compresses the model by 98.92%, does not suffer significant performance degradation compared to the model compressed by 95.87% using ITGC. In other words, NICKEL & DiME is 3.83 times more efficient. Additionally, NICKEL & DiME incurs only a 9.68% performance decrease compared to the full model at a compression rate of 74.96%.
[…]
StyleGAN2 FFHQ Full model [21] 14.90B - 4.02
ITGC [16] 3.73B 74.96% 5.27
NICKEL & DiME 4.42
GS [55] 1.38B 90.73% 10.26
CAGC [35] 10.06
GCC [32] 11.19
ITGC 10.02
DiME 8.39
NICKEL & DiME 7.43
CAGC 0.61B 95.87% 14.29
ITGC 14.01
DiME 11.25
NICKEL & DiME 10.45
CAGC 0.16B 98.92% 23.05
ITGC 22.23
DiME 18.32
NICKEL & DiME 15.93
LSUN-CAT Full model [21] 14.90B - 8.19
GS 1.38B 90.73% 17.11
CAGC 12.31
ITGC 12.06
DiME 11.59
NICKEL & DiME 10.80
(a) Comparison of FID
(b) Comparison of Precision
(c) Comparison of Recall
Figure 4: Performance comparison as a function of compression rates on StyleGAN2 for FFHQ. (a) indicates a function showing how FID varies with compression rates. NICKEL & DiME consistently outperforms other state-of-the-art compression methods at various compression rates. At a compression rate of 74.96%, NICKEL & DiME shows only 9.68% performance degradation compared to the full model, and the performance degradation due to increasing compression rates occurs less than other state-of-the-art compression methods. (b) indicates a function showing how Precision varies with compression rates. NICKEL & DiME shows comparable fidelity scores to other methods. (c) indicates a function showing how Recall varies with compression rates. NICKEL & DiME shows better preservation of diversity compared to other methods, even with higher compression rates.
4.1.2 Datasets.
We use FFHQ [20] and LSUN-CAT [59] datasets on StyleGAN2 [22], and CIFAR-10 [25] on SNGAN [36]. Resolution of FFHQ and LSUN-CAT is set to 256$\times$256.
4.1.3 Baselines.
We follow the architecture and training setups of StyleGAN2 [21], except for augmentation (not used). Our pruned generators are identical to CAGC [35]. The architecture and training setups for SNGAN follow Kang et al. [17]. We use the official code of StyleGAN2-ADA-PyTorch [^3] https://github.com/NVlabs/stylegan2-ada-pytorch and StudioGAN [^4] https://github.com/POSTECH-CVLab/PyTorch-StudioGAN. For clear comparison, we reproduce state-of-the-art GAN compression methods based on the StyleGAN2-ADA-PyTroch official code.
4.1.4 Evaluation metrics.
We use the Fréchet Inception Distance (FID) [11] metric to evaluate the performance of the GANs, measuring both the quality and diversity of the generated images. FID compares the features extracted by Inception V3555https://nvlabs-fi-cdn.nvidia.com/stylegan2-ada-pytorch/pretrained/metrics/inception-2015-12-05.pt between real and generated images by assuming Gaussian distribution. Precision & Recall are metrics for calculating the fidelity and diversity of data by computing the manifold of the data through K-Nearest Neighbors. We use VGG16666https://nvlabs-fi-cdn.nvidia.com/stylegan2-ada-pytorch/pretrained/metrics/vgg16.pt to compute Precision & Recall. In this paper, we use 50,000 samples and set the neighborhood size to 3. We note the average of the five best FID scores.
4.2 Results
We first compare the knowledge distillation performance of DiME, as described in Eq. 1, with state-of-the-art GAN compression methods[55, 35, 32, 16]. To distill the knowledge of $G^{T}$, DiME compares the outputs of $G^{T}$ and $G^{S}$ in the foundation embedding spaces (i.e., DINO, CLIP). As shown in Tab. 1 and Tab. 2, DiME outperforms the previous compression methods on various GAN architectures and datasets. Particularly, DiME improves FID scores by 1.63 compared to the state-of-the-art GAN compression methods in a setting where it reduces the FLOPS of StyleGAN2 for FFHQ by 11 times, with a compression rate of 90.73%. Our experimental results show that DiME is highly effective for knowledge distillation.
Additionally, to investigate the effectiveness of distillation considering the characteristics of GANs (i.e., NICKEL) beyond direct knowledge distillation (i.e., DiME), we combine NICKEL, as described in Eq. 2, and DiME. Tab. 1 and Tab. 2 show that NICKEL & DiME further improves the performance over that of DiME. In Tab. 2, our method achieves the FID score of 15.93 by compressing StyleGAN2 93-fold. This compares to ITGC, which attains the FID score of 14.01 with a 24-fold compression.
Table 3: Quantitative results of extremely compressed StyleGAN2. We compare the performance of various compression methods on StyleGAN2 for FFHQ at compression rate = 99.69%. Previous methods often suffer from severe performance degradation due to the imbalance between $G^{S}$ and $D^{S}$ when GAN is extremely compressed. On the other hand, NICKEL & DiME shows acceptable performance compared to other methods with high stability.
[…]
StyleGAN2 FFHQ Full model [21] 14.90B - 4.02
GS [55] 0.05B 99.69% 184.33
CAGC [35] 186.61
ITGC [16] 164.92
NICKEL & DiME 29.38
For in-depth investigations, we compare the FID, Precision, and Recall performance of compression methods at various compression rates. Fig. 4 indicates the FID, Precision, and Recall scores at each compression rate for StyleGAN2 on FFHQ dataset. In Fig. 4(a), NICKEL & DiME outperforms previous methods at all reported compression rates. Furthermore, at a compression rate of 74.96%, NICKEL & DiME shows only a 9.68% performance degradation (FID: 4.42), compared to the full model (FID:4.03). Remarkably, our method shows significant gaps with previous methods as compression rates increase, thanks to the improved stability. Fig. 4(b) shows precision scores, which indicate the fidelity of generated images. We observe that the precision scores of the compressed models are higher than those of the full model. While we observe a deterioration of precision scores with increasing compression rates, NICKEL & DiME maintains the precision scores comparable to the full model, even at high compression rates. Moreover, NICKEL & DiME shows precision scores comparable to the precision score of the full model up to a compression rate of 98.92%. Fig. 4(c) shows recall scores, indicating the diversity of the generated images. We observe that, unlike precision scores, the recall scores of the compressed models decrease significantly with increasing compression rates. Still, NICKEL & DiME maintains better diversity compared to the other compression methods. We provide performance comparison with recent various metrics in the supplementary for more comprehensive analysis.
(a) compression rate = 90.73%
(b) compression rate = 90.73%
Figure 5: Visualization of images generated by compressed StyleGAN2 on FFHQ and LSUN-CAT. (a) shows the visual quality of StyleGAN2 compressed by NICKEL & DiME on FFHQ at compression rate = 90.73%. (b) shows the visual quality of StyleGAN2 compressed by NICKEL & DiME on LSUN-CAT at compression rate = 90.73%.
As shown in Fig. 3, NICKEL & DiME mitigates the imbalance between $G^{S}$ and $D^{S}$ by considering $D^{S}$ during knowledge distillation. Fig. 3(a) and Fig. 3(b) respectively show the logits of $D^{S}$ for ITGC and NICKEL & DiME. In contrast to the ideal training of GAN where the logits of the discriminator should be close to 0, ITGC shows significant performance degradation due to the imbalance between $G^{S}$ and $D^{S}$ during training. Particularly, as the compression on generator intensifies, the imbalance between $G^{S}$ and $D^{S}$ becomes more pronounced. On the other hand, NICKEL & DiME alleviates this imbalance. Even at a compression rate of 98.92%, our method maintains a better equilibrium compared to ITGC’s at the compression rate of 90.73%. Fig. 3(c) shows the convergence of FID scores, indicating stable convergence of our method compared to the other alternatives. It is noteworthy that our method shows stable convergence even under extreme compression rates. As shown in Tab. 3, at an extreme compression rate of 99.69%, other methods fail to achieve stable learning due to the breakdown of Nash equilibrium between the highly compressed generator and discriminator. In contrast, our method not only shows stable convergence but also achieves reasonable performance, even with a 321-fold compression. Sec. 4.3 shows the visual quality of this scenario.
(a) compression rate = 98.92%
(b) compression rate = 99.69%
Figure 6: Visualization of extremely compressed StyleGAN2 on FFHQ. (a) indicates the visual quality of StyleGAN2 compressed by NICKEL & DiME at compression rate = 98.92%. (b) indicates the visual quality of StyleGAN2 compressed by NICKEL & DiME at compression rate = 99.69%. NICKEL & DiME shows acceptable visual quality even at extreme compression rates.
4.3 Visualization of a compression factor of 11, 92, and 321.
In Fig. 5, we show generated images for FFHQ and LSUN-CAT datasets using StyleGAN2 at a compression rate of 90.73%. Our method shows not only high visual quality but also the ability to generate diverse images. In Fig. 6, we visualize generated images at high compression rates. At a compression rate of 98.92%, our method shows visual quality that is not significantly degraded. Moreover, even at a compression rate of 99.69%, our method shows reasonable visual quality with diverse images.
4.4 Ablation Study
In this section, we study the power of kernel embedding spaces and the effect of components of our method.
CLIP and DINO embeddings. In Tab. 4, the CLIP embedding w/o global indicates using only CLIP as the embedding kernel for knowledge distillation. We observe significant challenges in achieving stable knowledge distillation when using only CLIP. On the other hand, when using only DINO as the embedding kernel, DINO embedding w/o global, we observe stable convergence and achieve the FID of 20.75. In addition, we observe that although the CLIP embedding space may pose challenges in achieving stable knowledge distillation, combining it with the DINO embedding space could lead to slight performance improvements.
Utilization of global features. The objective of KD is to match the population distributions between $G^{T}$ and $G^{S}$. However, due to the batch size, we can only match the sample distributions. Hence, a sampling error $\epsilon_{KD}$ may occur in the KD loss, which is bounded by the sum of the sampling errors of $G^{T}$ and $G^{S}$:
$\epsilon_{KD}<\epsilon_{teacher}+\epsilon_{student}$ (4)
Fortunately, unlike $G^{S}$, the distribution of $G^{T}$ is fixed. Therefore, by precomputing the statistics— referred to as global features —through infinite sampling, we can achieve an infinitesimal sampling error $\epsilon_{teacher}$. As shown in Tab. 4, we find that utilizing global features leads to performance enhancement. In fact, this resembles the MMD critic, which is a stable metric for learning the distribution. Santos et al. [48] and Yeo et al. [58] noted that pretrained neural networks can be considered as characteristic kernels, and reducing the discrepancy of the mean between extracted features can be seen as the MMD critic. In this vein, DiME can be considered to stably match distributions between two generators. In summary, we observe that combining NICKEL & DiME shows the best performance.
Table 4: Ablation study results in NICKEL & DiME.
[…]
DINO CLIP
CLIP embedding w/o global 152.15
DINO embedding w/o global 20.75
DINO embedding 19.60
DiME 18.32
NICKEL & DiME 15.93
5 Limitations
Our method shows excellent performance via distribution matching, yet it tends to focus on the fidelity of generated images. In fact, every method experiences significant degradation in recall performance, even at low compression rates (Figure 4). Furthermore, there still remains the imbalance between the generator and discriminator at extreme compression rates, which incurs significant performance degradation. Thus, it is an interesting research direction to develop methods that are capable of maintaining diversity and stability when compressing generative models at extreme compression rates.
6 Conclusion
In this paper, we propose Distribution Matching for Efficient compression (DiME) and Network Interactive Compression via Knowledge Exchange and Learning (NICKEL) that set a new standard of the performance in GAN compression. DiME matches the distributions between the teacher generator and student generator by using the maximum mean discrepancy (MMD) as a loss function. For better matching, we harness the power of the pretrained foundation model and use it as embedding kernels in MMD loss for knowledge distillation. DiME can compress StyleGAN2 with the FID of 4.02 by 20 times while maintaining reasonable performance with the FID of 11.25, achieving the state-of-the-art performance in all compression rates. NICKEL further enhances the performance by providing better feedback to the student generator from the discriminator. Combining these two, NICKEL & DiME successfully compresses StyleGAN2 by 92 times while maintaining the FID score of 15.93. Thanks to its enhanced stability, NICKEL & DiME allows us to compress StyleGAN2 by up to 99.69% (321 times smaller) while maintaining reasonable performance, which is not possible for existing methods.